OpenStudy (anonymous):

Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is tru

OpenStudy (campbell_st):

Part A  (b) the ratio of A to B or A:B is 2:1 B has measurement 1/2 of A Area ratio (1/2)^2 = 1/4 Volume ratio (1/2)^3 = 1/8 (c) Volume of A V = 6x4x2 = 48 Volume of B V = 3x2x1 = 6 Volume A x 1/8 = 48 x 1/8 = 6 the volume of B Surface Area